# Transit for All: Mapping Equitable Bike2Subway Connection using Region Representation Learning

### v1

@@ 4.3. Recommend New Bike Station Placement
To address gaps in equitable transit accessibility, we recommend locations for new bike-sharing stations in NYC based on wPTAL. Official Citi Bike guidelines for bike station placement (New York City Department of Transportation, 2013) focus on both geographic information and site-specific criteria. For example, they include minimum sidewalk widths (16 ft), curb-lane widths (8 ft), spacing of approximately 305 meters (1,000 ft) between stations, a 15-ft clearance from subway entrances, restrictions around fire hydrants, and bus stops, and several on-site observations, such as avoiding building entrances and construction area. However, these guidelines do not incorporate demand prediction (Berke et al., 2024; Mahajan and Argota Sánchez-Vaquerizo, 2024; Javid and Sadeghvaziri, 2023b), which is critical for ensuring that new stations address local community transit needs and maximize their actual utilization.
@@
  • Considering expected bike usage and transit accessibility: We select candidate station locations with wPTAL. Stations with higher wPTAL scores indicate areas where improved bike-sharing access would significantly enhance equitable transit accessibility and community-level mobility.
@@ 5.1.1. CooKIE + Spatial-MGAT input data:
  • Satellite imagery: We sample high-resolution (1-meter) aerial imagery from the 2021 National Agriculture Imagery Program (NAIP) (United States Department of Agriculture, 2021) within a 500-meter buffer around each cold-start station.
@@
5.1.2. wPTAL input data
To calculate wPTAL scores for cold-start stations, we combine predicted bike demand and PTAL score. To compute PTAL, we utilize subway entrances, pedestrian road networks, and transit schedules data. Specifically, we collect:
@@ 6. Results & Discussion
We first compare bike demand prediction at cold-start station using CooKIE + Spatial-MGAT versus raw features + Spatial-MGAT. Next, we visualize wPTAL distributions across NYC. We then perform zone-level analyses with census data across multiple scenarios and functional areas. Finally, we recommend future locations for potential BSS stations.
@@ 6.1. Demand Prediction at Cold-start Stations
We design an experiment that utilize multimodal representation learned by CooKIE and raw feature concatenation, which we feed these inputs to Spatial-MGAT. We evaluate demand prediction performance at cold-start stations between CooKIE and raw features concatenation approaches (Raw) across root mean squared error (RMSE), mean average error (MAE) and R-squared ($R^{2}$(Liang et al., 2023) in Table 1.
@@
These evaluations demonstrate that multimodal region representations learned via CooKIE outperform Raw + Spatial-MGAT, demonstrating their effectiveness for generating bike demand prediction in newly planned bike stations.
6.2. Spatial Distribution of wPTAL
We present the spatial distribution of wPTAL for newly planned bike-sharing stations from January 2022 to December 2024 (testing period) in Figure 4. Figure 4 shows Manhattan as having consistently high wPTAL scores, resulting from strong public transit infrastructure combined with high predicted ridership. In Queens, stations along major roads, particularly those aligned with subway lines, also exhibit high wPTAL values (shown in red), emphasizing the high public transport accessibility with the demand. In contrast, Brooklyn shows more variability with high wPTAL scores concentrated near major transit hubs but notably lower scores in neighborhood areas. The Bronx shows similarly mixed accessibility patterns with moderate wPTAL values primarily clustered around central roads and transit corridors.
To clearly illustrate the advantage of using wPTAL over PTAL alone, we closely examine three representative stations: Station A (Manhattan), Station B (Queens), and Station C (Brooklyn). Although their raw PTAL scores are similar (differing by at most 2.8 points), their wPTAL scores vary significantly. Specifically, Station A achieves a wPTAL of 1,436.3, substantially higher than Stations B (139.2) and C (133.7). This substantial difference indicates that Station A, which is located in a densely populated area with high predicted ridership, benefits from integrating ridership predictions into accessibility metrics. Thus, incorporating predicted ridership into wPTAL effectively captures realistic differences in transit accessibility and provides guidance for equitable and impactful station placements.
Figure 4. Spatial distribution of wPTAL for 699 new BSS stations during the test period (2022–2024). Red indicates higher wPTAL scores, identifying stations with both strong transit access and predicted ridership.
@@ 6.3. Zone-level Analysis across Scenarios
6.3.1. Zone-level Analysis: wPTAL per Major Ethnicity
@@
Figure 5. Spatial distribution of wPTAL for cold-start stations across major ethnic groups (White, Black, Asian, and Hispanic). Darker green indicates higher wPTAL scores, reflecting better public transportation accessibility. The map highlights distinct spatial accessibility patterns across ethnic groups.
@@
Figure 6. Spatial distribution of wPTAL for cold-start stations categorized by median income (high, medium-high, medium-low, low). Darker blue indicates higher public transport accessibility scores.
@@
We visualize the spatial distribution of wPTAL across four ethnic groups, White, Black, Asian, and Hispanic, in Figure 5. We observe that the spatial distribution of wPTAL distinctly varies across these ethnicity groups. The White population predominantly resides in central Manhattan and along the boundary between Brooklyn and Queens. These areas benefit from extensive subway infrastructure, resulting in notably high accessibility with a mean wPTAL of 7,262.4. In contrast, the Black population is largely concentrated in southern Brooklyn, a region with fewer subway lines and limited public transit coverage, resulting in significantly lower accessibility with a mean wPTAL of 459.1. Asian neighborhoods cluster prominently in Queens, especially around major transit hubs such as Flushing and Jackson Heights neighborhood. These areas benefit from multiple subway lines and bus services, which yields relatively high accessibility reflected by a mean wPTAL of 2,138.3. Hispanic neighborhoods, predominantly living in northern Brooklyn and eastern Queens, typically lie further from major transit routes, resulting in the lowest mean wPTAL among the four ethnic groups at 244.0. Nonetheless, some central Hispanic neighborhoods in Queens exhibit better transit accessibility than those at the borough’s edges or the Bronx, indicating localized variations in service distribution.
@@
Considerable variability within ethnic groups further highlights transit accessibility disparities: the standard deviation of wPTAL is notably high, such as 16,346.2 for White neighborhoods and 8,572.1 for Asian neighborhoods. These variations reflect disparities from not solely by ethnicity but also by neighborhood-specific urban planning and infrastructure. Analyzing wPTAL at the borough level further clarifies these patterns. Black neighborhoods shows consistently low accessibility, with mean wPTAL scores of 324.4 in Brooklyn and 615.2 in Queens. This large difference in wPTAL by boroughs indicates significant transit gaps and the need for improvements to enhance equity. On the other hand, Asian neighborhoods in Queens show higher accessibility, with a mean wPTAL score of 652.1, benefiting from proximity to robust transit infrastructure. Hispanic neighborhoods show lower borough-level accessibility, with mean wPTAL scores of 183.4 in Brooklyn and 235.3 in Queens, highlighting transit limitations due to their distant residential locations relative to major transit locations.
6.3.2. Zone-level Analysis: wPTAL per Median Income
@@
In Figure 6, we analyze the spatial distribution of wPTAL by dividing regions into four median-income groups based on quantiles: high, medium-high, medium-low, and low. The resulting spatial patterns distinctly reflect the economic dynamics across NYC. Stations in high median-income neighborhoods demonstrate the highest overall accessibility, with an average wPTAL of 9,503.7. These neighborhoods, predominantly located in central Manhattan, benefit significantly from extensive subway networks, frequent transit services, and robust infrastructure. High-income areas outside Manhattan, such as those in Brooklyn and Queens, still show relatively high accessibility but with notably lower mean wPTAL scores than in Manhattan: 1,130.9 in Brooklyn and 740.2 in Queens. Neighborhoods with medium-high income have an average wPTAL of 2,180.5 citywide, though substantial variations exist among boroughs. Medium-high income neighborhoods in Manhattan have exceptionally high accessibility, averaging 38,264.0, due to proximity to major transit hubs and dense subway coverage. In contrast, neighborhoods in the same income category in Brooklyn and Queens represent considerably lower accessibility than in Manhattan, averaging 337.1 and 246.0, respectively. Such disparities within the same income group highlight significant differences in transit accessibility across boroughs.
@@
Neighborhoods within the medium-low income category generally experience lower and more varied transit accessibility than high or medium-high categories, averaging a citywide wPTAL of 337.7. Within the medium-low income group, neighborhoods in Manhattan have comparatively higher accessibility than other boroughs, averaging 2,668.4, benefiting from nearby extensive transit infrastructure. However, medium-low income neighborhoods in the Bronx, Brooklyn, and Queens exhibit notably lower accessibility, averaging 172.6, 277.2, and 308.3, respectively. This large difference among boroughs underscores infrastructure disparities that significantly impact transit accessibility. Finally, neighborhoods with low-income groups consistently experience the lowest accessibility citywide, averaging a wPTAL of 425.6. Borough-level analysis reveals particularly low accessibility scores for low-income neighborhoods in the Bronx at 205.7, Brooklyn at 260.9, and Queens at 392.4. In contrast, Manhattan’s low-income neighborhoods exhibit notably higher accessibility than those in three boroughs, averaging 2,221.8. These borough-level differences emphasize severe inequities in transit accessibility, disproportionately affecting economically vulnerable communities, particularly those in the outer boroughs compared to Manhattan.
@@ 6.4. Zone-level Analysis across Functional Areas
Figure 7. Spatial distribution of wPTAL by dominant Points-of-Interest (POIs): Commercial, Education, Recreational, Residential, Social, and Transportation. A darker orange color indicates higher wPTAL scores.
@@
We analyze the spatial distribution of wPTAL across various POI categories: Commercial, Education, Recreational, Residential, Social, and Transportation zones (Figure 7). The spatial patterns of wPTAL distinctly align with both demographic (Section 5) and economic patterns (Section 6) in NYC. Commercial zones consistently exhibit high wPTAL, notably in Manhattan’s central business districts with an average wPTAL of 37,303.4. These areas benefit substantially from dense subway networks and frequent transit services. Brooklyn’s commercial zones also have relatively high accessibility, averaging 873.0, whereas Queens shows significantly lower scores, averaging 305.5. Transportation-related POIs generally display high wPTAL due to their connectivity and proximity to major transit hubs, such as transit terminals and ferry landings. Manhattan’s transportation zones indicate exceptionally high wPTAL, averaging 4,660.9. However, Queens and Brooklyn demonstrate markedly lower accessibility, averaging 409.5 and 579.0, respectively, reflecting variations in transit service frequency and density. Residential zones typically exhibit lower overall wPTAL, with borough-specific variations clearly visible. Residential areas in Manhattan show relatively high average accessibility scores at 1,790.4, benefiting from proximity to extensive transit infrastructure. In contrast, residential zones in Brooklyn, Queens, and the Bronx demonstrate significantly lower accessibility scores than Manhattan, averaging 195.3, 113.3, and 462.1, respectively. Particularly, low scores in southern Brooklyn and the Bronx align closely with areas predominantly by lower-income and minority communities.
@@
Educational zones present varied transit accessibility, heavily influenced by proximity to major transit. Manhattan educational institutions, including universities and large public schools, show extremely high average accessibility at 19,004.0. Conversely, educational facilities in Brooklyn (577.7), Queens (426.7), and especially the Bronx (211.8) display notably lower wPTAL scores, reflecting disparities in educational accessibility linked to transit infrastructure. Recreational POIs similarly display significant variability. Central Manhattan recreational facilities, benefiting from proximity to transit-rich landmarks and attractions, average 5,592.7 in accessibility. However, recreational zones in outer boroughs such as the Bronx (89.6), Queens (204.2), and Brooklyn (426.8) indicate substantially lower scores, particularly for remote locations such as beaches, golf courses, and gardens. Social service zones, including daycare centers, nursing homes, and shelters, show notably low wPTAL across boroughs. Manhattan’s social service facilities have high wPTAL averaging 23,274.0 due to central locations within transit-rich neighborhoods. Meanwhile, social service facilities in Queens (273.1), Brooklyn (284.7), and notably the Bronx (0.0), reflect significantly lower accessibility, highlighting critical transit inequities impacting vulnerable populations in these boroughs.
@@ 6.5. Future BSS Station Recommendations
We quantify equity improvements in wPTAL for future bike stations using the Gini index (Jiang et al., 2024; Allanson, 2014). The Gini index captures disparities in accessibility distribution, where a lower value indicates more equitable access across demographic (income) groups. The Gini index measures relative inequality among demographic groups by comparing group means to overall weighted accessibility. Specifically, the Gini index $G$ for a set of $M$ demographic groups is defined as:
@@
where $w_{i}$ and $w_{j}$ represent the number of stations in demographic groups $i$ and $j$ (e.g., income-based categories like high, medium-high), respectively, $m_{i}$ and $m_{j}$ are the mean wPTAL score within groups $i$ and $j$, $W$ is the total number of stations across all demographic groups, and $\mu$ is the population-weighted overall mean accessibility.
Figure 8. Comparison of Gini coefficients measuring equity in wPTAL by income and race, Citi Bike Plan vs. Transit for All, incrementally adding recommended bike stations. Lower Gini coefficients indicate better equity.
@@
Furthermore, we visualize the spatial distribution of our recommended stations in Figure 9. The recommended stations (blue) effectively target underserved neighborhoods in the Bronx, eastern Queens, and southern Brooklyn, complementing existing (gray) and previously planned (red) Citi Bike stations. Our wPTAL-based site selection demonstrates the strategic future station recommendation to close gaps in the current BSS network, enhancing transit equity by extending accessibility to communities previously underserved by bike-sharing infrastructure. This highlights the effectiveness of our TFA framework in improving transport accessibility equity, aligned with SDG 11.2.

### v2

@@ 4.3. Recommend New Bike Station Placement
To address gaps in equitable transit accessibility, we recommend locations for new bike-sharing stations in NYC based on wAL. An official Citi Bike guidelines for bike station placement (New York City Department of Transportation, 2013) focus on both geographic information and site-specific criteria. For example, they include minimum sidewalk widths (16 ft), curb-lane widths (8 ft), spacing of approximately 305 meters (1,000 ft) between stations, a 15-ft clearance from subway entrances, restrictions around fire hydrants, and bus stops, and several on-site observations, such as avoiding building entrances and construction area. However, these guidelines do not incorporate demand prediction (Berke et al., 2024; Mahajan and Argota Sánchez-Vaquerizo, 2024; Javid and Sadeghvaziri, 2023b), which is critical for ensuring that new stations address local community transit needs and maximize their actual utilization.
@@
  • Considering expected bike usage and transit accessibility: We select candidate station locations with wAL. Stations with higher wAL scores indicate areas where improved bike-sharing access would significantly enhance equitable transit accessibility and community-level mobility.
@@ 5.1.1. CooKIE + Spatial-MGAT input data:
  • Satellite imagery: We sample high-resolution (0.6-meter) aerial imagery from the 2021 National Agriculture Imagery Program (NAIP) (United States Department of Agriculture, 2021) within a 500-meter buffer around each cold-start station.
@@
5.1.2. wAL input data
To calculate wAL scores for cold-start stations, we combine predicted bike demand and AL score. To compute AL, we utilize subway entrances, pedestrian road networks, and transit schedules data. Specifically, we collect:
@@ 6. Results & Discussion
We first compare bike demand prediction at cold-start station using CooKIE + Spatial-MGAT versus raw features + Spatial-MGAT. Next, we visualize wAL distributions across NYC. We then perform zone-level analyses with census data across multiple scenarios and functional areas. Finally, we recommend future locations for potential BSS stations.
@@ 6.1. Demand Prediction at Cold-start Stations
We design an experiment that feeds multimodal representation learned by CooKIE and raw feature concatenation to Spatial-MGAT as inputs. We evaluate BSS demand prediction at cold-start stations during the test period (See Section 5.2) using root mean squared error (RMSE), mean average error (MAE), and R-squared ($R^{2}$(Liang et al., 2023) in Table 1.
@@
Figure 4. Spatial distribution of prediction errors for Raw + Spatial-MGAT and CooKIE + Spatial-MGAT (Ours). Larger errors appear in dense regions (shown with arrows) when using raw features, whereas CooKIE reduces these errors and yields more consistent predictions across the city.
Figure 4 shows the prediction errors across NYC. With raw feature concatenation, large errors cluster in high-demand areas such as the Bronx, Queens, and Brooklyn (yellow arrows), where bike demand is difficult to capture from simple features. In contrast, CooKIE reduces these errors, producing more stable predictions across diverse neighborhoods.
6.2. Spatial Distribution of wAL
We present the spatial distribution of wAL for newly planned BSS stations from January 2022 to December 2024 (testing period) in Figure 5. Figure 5 shows Manhattan as having consistently high wAL scores, resulting from strong public transit infrastructure combined with high predicted ridership. In Queens, stations along major roads, particularly those aligned with subway lines, also exhibit high wAL values (shown in red), emphasizing the high transport accessibility with the demand. In contrast, Brooklyn shows more variability with high wAL scores concentrated near major transit hubs but notably lower scores in neighborhood areas. The Bronx shows similarly mixed accessibility patterns with moderate wAL values primarily clustered around central roads and transit corridors.
To clearly illustrate the advantage of using wAL over AL alone, we closely examine three representative stations: Station A (Manhattan), Station B (Queens), and Station C (Brooklyn). Although their raw AL scores are similar (differing by at most 2.8 points), their wAL scores vary significantly. Specifically, Station A achieves a wAL of 1,436.3, substantially higher than Stations B (139.2) and C (133.7). This substantial difference indicates that Station A, which is located in a densely populated area with high predicted ridership, benefits from integrating ridership predictions into accessibility metrics. Thus, incorporating predicted ridership into wAL effectively captures realistic differences in transit accessibility and provides guidance for equitable and impactful station placements.
Figure 5. Spatial distribution of wAL for 699 new BSS stations during the test period (2022–2024). Red indicates higher wAL scores, identifying stations with both strong transit access and predicted ridership.
@@ 6.3. Zone-level Analysis across Scenarios
6.3.1. Zone-level Analysis: wAL per Major Ethnicity
@@
Figure 6. Spatial distribution of wAL for cold-start stations across major ethnic groups (White, Black, Asian, and Hispanic). Darker green indicates higher wAL scores, reflecting better public transportation accessibility. The map highlights distinct spatial accessibility patterns across ethnic groups.
@@
Figure 7. Spatial distribution of wAL for cold-start stations categorized by median income (high, medium-high, medium-low, low). Darker blue indicates higher accessibility scores.
@@
We visualize the spatial distribution of wAL across four ethnic groups, White, Black, Asian, and Hispanic, in Figure 6. We observe that the spatial distribution of wAL distinctly varies across these ethnicity groups. The White population predominantly resides in central Manhattan and along the boundary between Brooklyn and Queens. These areas benefit from extensive subway infrastructure, resulting in notably high accessibility with a mean wAL of 7,262.4. In contrast, the Black population is largely concentrated in southern Brooklyn, a region with fewer subway lines and limited public transit coverage, resulting in significantly lower accessibility with a mean wAL of 459.1. Asian neighborhoods cluster prominently in Queens, especially around major transit hubs such as Flushing and Jackson Heights neighborhood. These areas benefit from multiple subway lines and bus services, which yields relatively high accessibility reflected by a mean wAL of 2,138.3. Hispanic neighborhoods, predominantly living in northern Brooklyn and eastern Queens, typically lie further from major transit routes, resulting in the lowest mean wAL among the four ethnic groups at 244.0. Nonetheless, some central Hispanic neighborhoods in Queens exhibit better transit accessibility than those at the borough’s edges or the Bronx, indicating localized variations in service distribution.
@@
Considerable variability within ethnic groups further highlights transit accessibility disparities: the standard deviation of wAL is notably high, such as 16,346.2 for White neighborhoods and 8,572.1 for Asian neighborhoods. These variations reflect disparities from not solely by ethnicity but also by neighborhood-specific urban planning and infrastructure. Analyzing wAL at the borough level further clarifies these patterns. Black neighborhoods shows consistently low accessibility, with mean wAL scores of 324.4 in Brooklyn and 615.2 in Queens. This large difference in wAL by boroughs indicates significant transit gaps and the need for improvements to enhance equity. On the other hand, Asian neighborhoods in Queens show higher accessibility, with a mean wAL score of 652.1, benefiting from proximity to robust transit infrastructure. Hispanic neighborhoods show lower borough-level accessibility, with mean wAL scores of 183.4 in Brooklyn and 235.3 in Queens, highlighting transit limitations due to their distant residential locations relative to major transit locations.
6.3.2. Zone-level Analysis: wAL per Median Income
In Figure 7, we analyze the spatial distribution of wAL by dividing regions into four median-income groups based on quantiles: high, medium-high, medium-low, and low. The resulting spatial patterns distinctly reflect the economic dynamics across NYC.
@@ 6.3.2. Zone-level Analysis: wAL per Median Income
Stations in high median-income neighborhoods demonstrate the highest overall accessibility, with an average wAL of 9,503.7. These neighborhoods, predominantly located in central Manhattan, benefit significantly from extensive subway networks, frequent transit services, and robust infrastructure. High-income areas outside Manhattan, such as those in Brooklyn and Queens, still show relatively high accessibility but with notably lower mean wAL scores than in Manhattan: 1,130.9 in Brooklyn and 740.2 in Queens.
Neighborhoods with medium-high income have an average wAL of 2,180.5 citywide, though substantial variations exist among boroughs. Medium-high income neighborhoods in Manhattan have exceptionally high accessibility, averaging 38,264.0, due to proximity to major transit hubs and dense subway coverage. In contrast, neighborhoods in the same income category in Brooklyn and Queens represent considerably lower accessibility than in Manhattan, averaging 337.1 and 246.0, respectively. Such disparities within the same income group highlight significant differences in transit accessibility across boroughs.
@@
Neighborhoods within the medium-low income category generally experience lower and more varied transit accessibility than high or medium-high categories, averaging a citywide wAL of 337.7. Within the medium-low income group, neighborhoods in Manhattan have comparatively higher accessibility than other boroughs, averaging 2,668.4, benefiting from nearby extensive transit infrastructure. However, medium-low income neighborhoods in the Bronx, Brooklyn, and Queens exhibit notably lower accessibility, averaging 172.6, 277.2, and 308.3, respectively. This large difference among boroughs underscores infrastructure disparities that significantly impact transit accessibility.
Finally, neighborhoods with low-income groups consistently experience the lowest accessibility citywide, averaging a wAL of 425.6. Borough-level analysis reveals particularly low accessibility scores for low-income neighborhoods in the Bronx at 205.7, Brooklyn at 260.9, and Queens at 392.4. In contrast, Manhattan’s low-income neighborhoods exhibit notably higher accessibility than those in three boroughs, averaging 2,221.8. These borough-level differences emphasize severe inequities in transit accessibility, disproportionately affecting economically vulnerable communities, particularly those in the outer boroughs compared to Manhattan.
@@ 6.4. Zone-level Analysis across Functional Areas
Figure 8. Spatial distribution of wAL by dominant Points-of-Interest (POIs): Commercial, Education, Recreational, Residential, Social, and Transportation. A darker orange color indicates higher wAL scores.
We analyze the spatial distribution of wAL across various POI categories: Commercial, Education, Recreational, Residential, Social, and Transportation zones, shown in Figure 8. The spatial patterns of wAL distinctly align with both demographic (Section 6) and economic patterns (Section 7) in NYC.
@@
Commercial zones consistently exhibit high wAL, notably in Manhattan’s central business districts with an average wAL of 37,303.4. These areas benefit substantially from dense subway networks and frequent transit services. Brooklyn’s commercial zones also have relatively high accessibility, averaging 873.0, whereas Queens shows significantly lower scores, averaging 305.5. Transportation-related POIs generally display high wAL due to their connectivity and proximity to major transit hubs, such as transit terminals and ferry landings.
Manhattan’s transportation zones indicate exceptionally high wAL, averaging 4,660.9. However, Queens and Brooklyn demonstrate markedly lower accessibility, averaging 409.5 and 579.0, respectively, reflecting variations in transit service frequency and density. Residential zones typically exhibit lower overall wAL with borough-specific variations clearly visible. Residential areas in Manhattan show relatively high average accessibility scores at 1,790.4, benefiting from proximity to extensive transit infrastructure. In contrast, residential zones in Brooklyn, Queens, and the Bronx demonstrate significantly lower accessibility scores than Manhattan, averaging 195.3, 113.3, and 462.1, respectively. Particularly, low scores in southern Brooklyn and the Bronx align closely with areas predominantly by lower-income and minority communities.
@@
Educational zones present varied transit accessibility, heavily influenced by proximity to major transit. Manhattan educational institutions, including universities and large public schools, show extremely high average accessibility at 19,004.0. Conversely, educational facilities in Brooklyn (577.7), Queens (426.7), and especially the Bronx (211.8) display notably lower wAL scores, reflecting disparities in educational accessibility linked to transit infrastructure.
Recreational POIs similarly display significant variability. Central Manhattan recreational facilities, benefiting from proximity to transit-rich landmarks and attractions, average 5,592.7 in accessibility. However, recreational zones in outer boroughs such as the Bronx (89.6), Queens (204.2), and Brooklyn (426.8) indicate substantially lower scores, particularly for remote locations such as beaches or golf courses. Social service zones, including daycare centers, nursing homes, and shelters, show notably low wAL across boroughs. Manhattan’s social service facilities show high wAL averaging 23,274.0 due to central locations within transit-rich neighborhoods. Meanwhile, social service facilities in Queens (273.1) and Brooklyn (284.7) reflect lower accessibility, highlighting critical transit inequities impacting vulnerable populations.
@@ 6.5. Future BSS Station Recommendations
We quantify equity improvements in wAL for future bike stations using the Gini index (Jiang et al., 2024; Allanson, 2014). The Gini index captures disparities in accessibility distribution, where a lower value indicates more equitable access across demographic (income) groups. The Gini index ($G$) measures relative inequality among demographic groups by comparing group means to overall weighted accessibility. Specifically, the $G$ for a set of $M$ demographic groups is defined as:
@@
where $w_{i}$ and $w_{j}$ represent the number of stations in demographic groups $i$ and $j$ (e.g., income-based categories like high, low), respectively, $m_{i}$ and $m_{j}$ are the mean wAL score within groups $i$ and $j$, $W$ is the total number of stations across all demographic groups, and $\mu$ is the population-weighted overall mean accessibility.
Figure 9. Comparison of Gini coefficients measuring equity in wAL by income and race, Citi Bike Plan vs. Transit for All, incrementally adding recommended bike stations. Lower Gini coefficients indicate better equity.
@@
Furthermore, we visualize the spatial distribution of our recommended stations in Figure 10. The recommended stations (blue) effectively target underserved neighborhoods in the Bronx, eastern Queens, and southern Brooklyn, complementing existing (gray) and previously planned (red) Citi Bike stations. Our wAL-based site selection demonstrates the strategic future station recommendation to close gaps in the current BSS network, enhancing transit equity by extending accessibility to communities previously underserved by bike-sharing infrastructure. This highlights the effectiveness of our TFA framework in improving transport accessibility equity, aligned with SDG 11.2.